OpenStudy (anonymous):

A hiker starts at his camp and moves the following distances while exploring his surroundings: 70.0 m north, 2.60 x 10^2m east, 110 m at an angle 30.0° north of east, and 1.50 x 10^2m south. (a) Find his resultant displacement from camp. (Take east as the positive x-direction and north as the positive y-direction.) magnitude ____ m direction ____ ° south of east (b) Would changes in the order in which the hiker makes the given displacements alter his final position? Explain.

OpenStudy (anonymous):

It's not clear from the problem as written what "2.60" and "1.50" are. Can shed some light on that?

OpenStudy (anonymous):

Sorry! It's corrected, thanks for letting me know!

OpenStudy (anonymous):

So if each of those moves is a vector, we need to assign a coordinate system and signs to the vectors. So let's say that north is positive y, and south is negative y. East is positive x, and west is negative x.

OpenStudy (anonymous):

Hopefully you know that vectors can have components. In this case of our coordinate system, a vector can have an x component and a y component. Note however that any vector that is parallel to the x-axis will only have an x-component. Likewise, any vector parallel to the y-axis will have only y component.

OpenStudy (anonymous):

When we add vectors, we add like components. So if one vector is u=<v,w> and another is r=<s,t>, u+r=<(v+s),(w+t)>.

OpenStudy (anonymous):

Three of your given vectors are parallel to the axes of the coordinate system. One isn't, and for that one you need to find its components.

OpenStudy (anonymous):

The x component of a vector is given by:\[v _{x}=||v||\cos \left( \theta \right)\]and the y component is given by:\[v _{y}=||v||\sin \left( \theta \right)\]This means that the vector v can be written as:\[v=<v _{x},v _{y}>\]

OpenStudy (anonymous):

Are beginning to see how to do the problem?

OpenStudy (mrnood):